# SKYSURF VI: The Impact of Thermal Variations of HST on Background Light Estimates

### v1

@@ 2 HST Temperature Data
To conduct this analysis, we accessed all publicly available archival WFC3 data from the Barbara A. Mikulski Archive for Space Telescopes (MAST) from June 2009 through December 2022. Specifically, we obtained all “spt”, “jit”, and “ima” fits files associated with WFC3/IR exposures. We chose to specifically focus on WFC3/IR exposures because the effect of thermal noise on shorter wavelengths is minimal (O’Brien et al., 2023). From the “spt” files, we obtained temperature data for the infrared focal plane array, optical bench near the IR fold mirror, infrared fold mirror, optical bench near the IR M1 mount, optical bench on the top cover, optical bench cold plate, detector radiator on the IR side by the condenser, detector radiator on IR side away from saddle, optical bench cold plate condenser saddle, and IR detector baseplate evaporator saddle. Particular attention was paid to the pick-off arm temperature ($T_{\rm POM}$) due to its increased sensitivity to changes in temperature and to sensors near the optical bench, because of its proximity to the filter wheel, where most of the thermal effect is expected to come from. Table 1 shows the default temperature values (referred to as $T_{\rm ref}$ throughout this work) of the HST optical components and the thermal background they contribute.
@@ 3 Short Term Temperature Changes
Figure 2: Time since occultation (seconds) versus average $T_{\rm POM}$ (in degrees Celsius) with orbital phase in 45 degree bins. For any time value greater than 6000 seconds, a value of 6000s is assigned. Data was obtained for each month of January only for the years 2015 through 2023.
Figure 3: Orbital phase (degrees) versus average $T_{\rm POM}$ (degrees Celsius). The orbital phase is binned over 50 degrees, and data are grouped by time since occultation in 1000s bins. Data was obtained for each month of January for 2015 through 2023.
Figure 3 shows how $T_{\rm POM}$ is impacted by orbital phase. $T_{\rm POM}$ changes with orbital phase by less than 1C. Temperature increases where orbital phase goes from 0 to 100 degrees, decreases from 100 to 250 degrees, then increases again. Overall, we find that $T_{\rm POM}$ is stable over the course of its orbit and is therefore not affected by earthshine. At orbital phases smaller than 270°, a negative correlation was observed, with correlation coefficients ranging from --0.96 to --0.10. At orbital phases larger than 270°, the correlation coefficients are positive. From 270 to 315°, the correlation coefficient is --0.89, and from 315 to 360°, the correlation coefficient is 0.65°. However, the p-values for all ranges of significance are greater than 0.05, with the exception of the 45–90° group, whose correlation coefficient has a p-value of 0.04. Therefore, we cannot consider these correlations to be significant. While there is no clear trend with time since occultation, there is a distinction between $T_{\rm POM}$ at different orbital phases.
@@ 4 Long Term Temperature Changes
Figure 7: Model of the deviation in $T_{\rm POM}$ from the average yearly temperature as a function of the Earth-Sun distance.
@@ 5 Empirical Thermal Background Constraints
where $Z(\alpha,\delta)$ represents Zodiacal emission at a given Right Ascension ($\alpha$) and Declination ($\delta$), $DGL$ represents diffuse emission, $IGL$ represents integrated galaxy light below HST’s detection limit, and $T$ represents the default thermal emission based on the thermal model of Carleton et al. (2022), corrected for the POM temperature variations described in Sec. 4. The Zodiacal spectrum is modeled as a reddened Solar spectrum as in
@@
where $S_{\odot}(\lambda)$ is the solar spectrum and $R$ is the reddening implied by the 1.25 $\mu$m and $2$ $\mu$m COBE/DIRBE data. The HST measurements have a much higher spatial resolution, such that deviations from that predicted reddening could certainly be apparent in our data. For this reason — and to allow our fit to be flexible — we allow the reddening to be modified by parameter $b$, as well as the Zodiacal normalization $a$. This has the flexibility to be somewhat model-independent and fit most astrophysical backgrounds, while having some physical basis and incorporating the (minor) known background variations within the visit.
For each visit, we fit (by minimizing $\chi^{2}$ between the model and observations) for $a$ and $b$, including only the F105W, F110W, F125W, and F140W data (for which the thermal background is comparatively small). We then extrapolate the model to the F160W observations. The difference between the measured and extrapolated backgrounds is taken to be the thermal background.
To compare the direct temperature measurements in Sec. 4 with these empirical thermal measurements, we model HST’s thermal background as a function of direct temperature measurements using pysynphot (STScI Development Team, 2013). This was done following the method of Carleton et al. (2022). Briefly, thermal emission is built up following the WFC3 optical path. For each optical component, the thermal emission is modeled with a modified blackbody with effective temperature corresponding to the temperature noted in the pysynphot reference files (with the keyword DEFT), modified by the wavelength-dependent emissivity noted in the same reference files. This spectrum was passed through the transmission of the remaining optical components until the light arrives at the WFC3/IR detector itself. For the default component temperatures, the predicted thermal background matches the thermal background listed in the instrument handbook, and we are able to modify the predicted thermal background level based on modified optical component temperatures. The code is hosted on github here: https://github.com/timcarleton/wfc3_thermal. Given the results of Sec. 3 and 4, we expect the temperature of the WFC3 optics to be constant (to within $<1\hbox{\,${}^{\circ}$C}$). The temperature of the external optics (the primary and secondary mirror), however, may vary more in principle.
@@ 6 Discussion of our Results
Figure 12: Top: Sky surface brightness measurements from this work for maximum and minimum $(T_{POM})$ and M1 temperature measurements, compared to those from Bernstein et al. (2002), Matsuura et al. (2017), and O’Brien et al. (2023), and the zodiacal light models from Aldering (2001) and Kelsall et al. (1998). The vertical line in the top left of the plot represents the average size of the error bars for these measurements. Errors are assumed to be approximately 3%, per Windhorst et al. (2022). Bottom: A zoomed-in version of the top plot to highlight the differences between instrument components and maximums and minimums.

### v2

@@ 2 HST Temperature Data
To conduct this analysis, we accessed all publicly available archival WFC3 data from the Barbara A. Mikulski Archive for Space Telescopes (MAST) from June 2009 through December 2022. Specifically, we obtained all “spt”, “jit”, and “ima” fits files associated with WFC3/IR exposures. To obtain these data, we filtered for all data types and selected only the WFC3 instrument. We also added the column ”aperture” and set the condition to ”IR”. Then, we filtered by date to select all data between June 2009 and December 2022. We chose to specifically focus on WFC3/IR exposures because the effect of thermal noise on shorter wavelengths is minimal (O’Brien et al., 2023). From the “spt” files, we obtained temperature data for the infrared focal plane array, optical bench near the IR fold mirror, infrared fold mirror, optical bench near the IR M1 mount, optical bench on the top cover, optical bench cold plate, detector radiator on the IR side by the condenser, detector radiator on IR side away from saddle, optical bench cold plate condenser saddle, and IR detector baseplate evaporator saddle. Particular attention was paid to the pick-off arm temperature ($T_{\rm POM}$) due to its increased sensitivity to changes in temperature and to sensors near the optical bench, because of its proximity to the filter wheel, where most of the thermal effect is expected to come from. Table 1 shows the default temperature values (referred to as $T_{\rm ref}$ throughout this work) of the HST optical components and the thermal background they contribute. Table 1 shows the default temperature values (referred to as $T_{\rm ref}$ throughout this work, and taken from published reference tables [^1] https://hst-crds.stsci.edu/) of the HST optical components and the thermal background they contribute, as determined from the Carleton et al. (2022) thermal model.
@@ 3 Short Term Temperature Changes
Figure 2: Time since occultation (seconds) versus average $T_{\rm POM}$ (in degrees Celsius) with orbital phase in 45 degree bins. For any time value greater than 6000 seconds, a value of 6000s is assigned. Data was obtained for each month of January only for the years 2015 through 2023. Errors are calculated as the standard deviation of the residuals in Figure 7.
Figure 3: Orbital phase (degrees) versus average $T_{\rm POM}$ (degrees Celsius). The orbital phase is binned over 50 degrees, and data are grouped by time since occultation in 1000s bins. Data was obtained for each month of January for 2015 through 2023. Errors are calculated as the standard deviation of the residuals in Figure 7.
Figure 3 shows how $T_{\rm POM}$ is impacted by orbital phase. $T_{\rm POM}$ changes with orbital phase by less than 1C. Temperature increases where orbital phase goes from 0 to 100 degrees, decreases from 100 to 250 degrees, then increases again. Overall, we find that $T_{\rm POM}$ is stable over the course of its orbit and is therefore not affected by earthshine. At orbital phases smaller than 270°, a negative correlation was observed, with correlation coefficients ranging from --0.96 to --0.10. At orbital phases larger than 270°, the correlation coefficients are positive. From 270 to 315°, the correlation coefficient is --0.89, and from 315 to 360°, the correlation coefficient is 0.65. However, the p-values for all ranges of significance are greater than 0.05, with the exception of the 45–90° group, whose correlation coefficient has a p-value of 0.04. Therefore, we cannot consider these correlations to be significant. While there is no clear trend with time since occultation, there is a distinction between $T_{\rm POM}$ at different orbital phases.
@@ 4 Long Term Temperature Changes
Figure 7: Model of the deviation in $T_{\rm POM}$ from the average yearly temperature as a function of the Earth-Sun distance. Error bars are calculated as the standard deviation of the residual.
@@ 5 Empirical Thermal Background Constraints
where $Z(\alpha,\delta)$ represents Zodiacal emission at a given Right Ascension ($\alpha$) and Declination ($\delta$), $DGL$ represents diffuse Galactic emission, $IGL$ represents integrated galaxy light below HST’s detection limit, and $T$ represents the default thermal emission based on the thermal model of Carleton et al. (2022), corrected for the POM temperature variations described in Sec. 4. Any additional diffuse component is accounted for by fitting for the overall normalization of the Zodiacal emission ($a$) as described below. The Zodiacal spectrum is modeled as a reddened Solar spectrum as in
@@
where $S_{\odot}(\lambda)$ is the solar spectrum and $R$ is the reddening implied by the 1.25 $\mu$m and $2.2$ $\mu$m COBE/DIRBE data (i.e. we found the $R$ value that best fit the ratio of the 1.25 $\mu$m and $2.2$ $\mu$m colors in the Kelsall et al. (1998) model). The $R$ values are mostly dependent on Sun Angle (with only a mild dependence on Ecliptic Latitude), ranging from $0.9$ at Sun Angles less than 100 deg to $0.2$ at Sun Angles greater than 140 deg. The HST measurements have a much higher spatial resolution, such that deviations from that predicted reddening could certainly be apparent in our data. For this reason — and to allow our fit to be flexible — we allow the reddening to be modified by parameter $b$, as well as the Zodiacal normalization $a$. This has the flexibility to be somewhat model-independent and fit most astrophysical backgrounds, while having some physical basis and incorporating the (minor) known background variations within the visit.
For each visit, we fit (by minimizing $\chi^{2}$ between the model and observations) for $a$ and $b$, including only the F105W, F110W, F125W, and F140W data (for which the thermal background is comparatively small). We find $a$ values ranging from $0.97$ to $1.08$ and $b$ values ranging from $-0.03$ to $1.7$ (both referring to the $10-90$th percentile range). This suggests that the deviations from the fiducial Kelsall et al. (1998) model are small. The assumed reddening is modified slightly more, but this may be expected given that only $1.25$ $\mu$m and $2.2$ $\mu$m bands were used to compute the expected reddening. Our results do not change if $R$ is taken to be a constant, suggesting that the HST data on its own is able to determine the reddening independently. The fit $a$ values increase slightly at high ecliptic latitudes and decrease slightly at high Sun Angles. The fit $b$ values show no strong correlation with sky position. We then extrapolate the model to the F160W observations. The difference between the measured and extrapolated backgrounds is taken to be the thermal background.
To compare the direct temperature measurements in Sec. 4 with these empirical thermal measurements, we model HST’s thermal background as a function of direct temperature measurements using pysynphot (STScI Development Team, 2013). This was done following the method of Carleton et al. (2022). Briefly, thermal emission is built up following the WFC3 optical path. For each optical component, the thermal emission is modeled with a modified blackbody with effective temperature corresponding to the temperature noted in the pysynphot reference files (with the keyword DEFT), modified by the wavelength-dependent emissivity noted in the same reference files. This spectrum was passed through the transmission of the remaining optical components until the light arrives at the WFC3/IR detector itself. For the default component temperatures, the predicted thermal background matches the thermal background listed in the instrument handbook, and we are able to modify the predicted thermal background level based on modified optical component temperatures. The code is hosted on github here: https://github.com/timcarleton/wfc3_thermal, and a version frozen at the time of submission is here: https://zenodo.org/records/13891457. Given the results of Sec. 3 and 4, we expect the temperature of the WFC3 optics to be constant (to within $<1\hbox{\,${}^{\circ}$C}$). The temperature of the external optics (the primary and secondary mirror), however, may vary more in principle.
@@ 6 Discussion of our Results
Figure 11: Top: Sky surface brightness measurements from this work for maximum and minimum $(T_{POM})$ and M1 temperature measurements, compared to those from Bernstein et al. (2002), Matsuura et al. (2017), and O’Brien et al. (2023), and the zodiacal light models from Aldering (2001) and Kelsall et al. (1998). The Bernstein et al. (2002) results are at an ecliptic longitude of $35.5$°and ecliptic latitude of $-35.5$°, and a galactic longitude of $206.6$°and a galactic latitude of $-59.8$°. The Matsuura et al. (2017) results are at an ecliptic longitude of $342.70$°and ecliptic latitude of $89.64$°, and a galactic longitude of $96.26$°and a galactic latitude of $29.46$°. Given that the SKYSURF points are normalized to the ecliptic pole, slight differences between previous measurements and SKYSURF measurements may be attributable to slightly different ecliptic and galactic positions. The vertical line in the top left of the plot represents the average size of the error bars for these measurements. Errors are assumed to be approximately 3%, per Windhorst et al. (2022). Bottom: A zoomed-in version of the top plot to highlight the differences between instrument components and maximums and minimums.